# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v1

@@ 2.4. Deployment Stage
2.4.1. Transfer Learning with Fine-tuning
@@
The typical steps of transfer learning involve first training a Wi-Fi sensing model on source domain data, followed by updating the model using target domain samples. This update can be applied to the task head, feature extractor, or the entire model. For example, XRF55 (Wang et al., 2024b), BullyDedect (Lan et al., 2024), and WiLISensing (Ding et al., 2020b) train models with source domain data and fine-tune the task head using several samples from the target domain, while freezing other parts of the model, enabling few-shot cross-domain activity recognition. DADA-AD (Chen et al., 2021) fine-tune the deep layers of the model. Besides, Brinke & Meratnia (Brinke and Meratnia, 2019b) experimentally compare the performance of fine-tuning the task head versus the entire network, finding that fine-tuning the whole network with 20% of target domain data yields the best results. In addition, in some cases, the target domain classification task differs from the source domain, such as in the number of action categories. To address this, WiTransfer (Fang et al., 2020) replaces the classification head and fine-tunes the new head using target domain data. FewSense (Yin et al., 2022), on the other hand, discards the classification head and fine-tunes the feature extractor using a cosine similarity loss. This loss pulls features of the same class closer while pushing features of different classes apart. During inference, it employs distance matching for classification, enabling the recognition of novel classes.
@@
Another type of transfer learning consists of two stages: the meta-training stage and the meta-testing stage. In the meta-training stage, a rich labeled dataset from the source domain is used to train an embedding model or feature extractor. In the meta-testing stage, the embedding model is frozen, and a small number of samples from the target domain are input for feature extraction. These features are then used to fine-tune a task head, allowing the model to adapt to the target domain with limited data. Several typical works employ this framework, such as ResMon (Zheng et al., 2023), which uses BayesCNN as the embedding model for domain-adaptive respiration state monitoring. RoGER (Bi et al., 2024) leverages CNN as the feature extractor to achieve domain-robust gesture recognition. Hou et al. (Hou et al., 2022a) and DASECount (Hou et al., 2022b) both use CNN networks as feature extractors to achieve cross-domain crowd counting.
2.4.2. Federated Learning
@@
WiFederated (Hernandez and Bulut, 2021), pFedBKD (Geng et al., 2023), CARING (Li et al., 2023), and Qi et al. (Qi et al., 2023) utilize federated learning frameworks to achieve efficient model deployment. Their core components include: (1) numerous clients with data collected from diverse environments, (2) a central server that trains a global model, which is expected to possess cross-environment generalization capabilities, and (3) the distribution of the global model to client devices, where it is fine-tuned using a small amount of local data. By leveraging these federated learning principles, Wi-Fi sensing models can be effectively deployed in unseen client environments.
@@
Some works adopt frameworks similar to federated learning but use different terminology. For instance, Co-WiSensing (Zhang et al., 2023d) proposes a cloud-edge collaborative framework. In this approach, a cloud server trains a cloud Wi-Fi sensing model using data from various edge devices. The higher layers of the trained model are pruned, and the lower layers are distributed to the edge devices. At the edge, the pruned model is extended with a task-specific head, which is fine-tuned using the edge user’s local data to generate a Wi-Fi sensing model tailored for the edge environment. Zhang et al. (Zhang et al., 2021b) also employs a central server and adversarial learning to extract shared attributes across different environments. The activity-oriented attributes are then transferred to the local server, enabling cross-environment Wi-Fi human activity recognition. AdaWiFi (Zheng et al., 2024) proposes a collective sensing framework, where a master device collects data using multiple sensors and trains multiple encoders collaboratively, which are then aggregated into a unified model. In the end-user environment, the model is collaboratively tuned with a small amount of labeled data. Zhang et al. (Zhang et al., 2023f) propose a local–global modeling method for indoor localization. The approach involves clustering Wi-Fi fingerprints of a large environment into multiple sub-areas, training local localization models for each sub-area, and then aggregating these local models into a global model. The authors demonstrate that this method achieves superior cross-environment generalization.
2.4.3. Continuous Learning
@@
After model deployment, it is desirable to continuously update the model to adapt to new environments or tasks. Some works achieve this with online learning (Shalev-Shwartz et al., 2012), a machine learning paradigm where the model learns and updates incrementally as new data becomes available. For instance, FewCS (Yang et al., 2023b) collects a small amount of online data to fine-tune model parameters in new environments, enabling few-shot cross-environment human action recognition. Similarly, WiDrive (Bai et al., 2019) utilizes online model adaptation and EM algorithm (Dempster et al., 1977) to update the model when in-car driver activities are misclassified, improving recognition for diverse vehicles and drivers. Unlike WiDrive’s automatic error detection, the User-in-the-loop approach allows users to report misclassifications or annotate a small number of samples, which are then used to update the model. M-WiFi (Soltanaghaei et al., 2020) specifically asks users to annotate critical time periods to fine-tune the trained model. To minimize the effort required from users, RISE (Zhai et al., 2021) combines probability and statistical assessments with anomaly detection to identify samples that are likely to be misclassified, reducing the need for users to report every error.
@@
Class incremental learning and domain incremental learning are also applied to enable Wi-Fi human sensing models to adapt to new activities or environments. For example, CCS (Fu et al., 2024) uses a data replay-based incremental learning approach (Zhao et al., 2020), where the model is updated to sense new human activity categories using representative samples from previous stages and samples from the current stage. Similarly, CSI-ARIL (Zhang et al., 2023c) treats the environment as a domain, updating the model with representative samples from both old and current environments. Both CCS and CSI-ARIL ensure that the model does not suffer from catastrophic forgetting (Kirkpatrick et al., 2017) when adapting to new classes or domains, preserving its ability to sense old tasks or domains.
3. Dataset
Table 1. Open-sourced Dataset in Wi-Fi Human Sensing.
ID dataset (#E, #p, #A) Device setting Access
@@ 3. Dataset
1 XRF55 (Wang et al., 2024b) 4 env., 39 part., 55 act. 1Tx, 3Rx, 20MB, 5GHz https://aiotgroup.github.io/XRF55/
2 Widar 3.0 (Zheng et al., 2019) 4 environment, 39 participants, 55 actions 1Tx, 3Rx, 20MB, 5GHz https://tns.thss.tsinghua.edu.cn/widar3.0/
3 ARIL (Wang et al., 2019a) 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/geekfeiw/ARIL
4 BullyDetect (Lan et al., 2024) 8 environments, 20 paired, 7 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/aiotgroup/WiFi-BullyDetect
5 HTHI (Alazrai et al., 2020) 1 environments, 40 paired, 12 actions 1Tx, 1Rx, 20MB, 5GHz https://data.mendeley.com/datasets/3dhn4xnjxw/1
6 WiAR (Guo et al., 2019) 5 environments, 10 participants, 16 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/linteresa/WiAR
7 Meneghello et.al. 2023  (Meneghello et al., 2023) 7 environments, 13 participants, 8 actions 1Tx, 1Rx, 20MB, 5GHz https://paperswithcode.com/paper/a-csi-dataset-for-wireless-human-sensing-on
8 Demrozi et.al. 2017 (Demrozi et al., 2023) 2 environments, 6 participants, 5 actions 1Tx, 1Rx, 20MB, 5GHz https://www.techrxiv.org/doi/full/10.36227/techrxiv.22284421.v1
9 Yousef et.al. 2017 (Yousefi et al., 2017) 1 environments, 6 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/ermongroup/Wifi_Activity_Recognition
10 WiSDA (Jiao et al., 2024) 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://www.researchgate.net/publication/385918549_WiSDA_Subdomain_Adaptation_Human_Activity_Recognition_Method_Using_Wi-Fi_Signals
11 WiNDR (Qin et al., 2024) 1 environments, 3 participants, 5 actions 3Tx, 3Rx, 20MB, 5GHz https://gitlab.com/yuxiqin/direction-independent
12 CSI-HAR-Dataset  (Moshiri et al., 2021) 3 environments, 6 participants, 5 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/parisafm/CSI-HAR-Dataset
13 WiMANS (Huang et al., 2024) 3 environments, 6 participants, 9 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/huangshk/WiMANS
14 ImgFi (Zhang and Jiao, 2023) 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://ieeexplore.ieee.org/document/10190332
15 ReWis (Bahadori et al., 2022) 3 environments, 2 participants, 3 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/niloobah/ReWiS
16 SAR (Brinke and Meratnia, 2019a, c) 1 environments, 9 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://data.4tu.nl/articles/_/12692816/1
17 DeepSeg (Xiao et al., 2020) 1 environments, 5 participants, 10 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/ChunjingXiao/DeepSeg
18 OPERAnet (Bocus et al., 2022) 2 environments, ??? participants, 7 actions 1Tx, 1Rx, 20MB, 5GHz https://springernature.figshare.com/collections/A_Comprehensive_Multimodal_Activity_Recognition_Dataset_Acquired_from_Radio_Frequency_and_Vision-Based_Sensors/5551209/1
19 NTU-Fi HAR (Yang et al., 2022a) 1 environments, 20 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/xyanchen/WiFi-CSI-Sensing-Benchmark
20 WiGuesture (Zhao et al., 2024b) 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/RS2002/CSI-BERT
21 RF-Net (Ding et al., 2020a) 6 environments, 11 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/di0002ya/RFNet
22 RISE (Zhai et al., 2021) 2 environments, 6 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/jiaojiao1234/RISE/
23 SignFi (Ma et al., 2018) 2 environments, 5 participants, 276 actions 3Tx, 1Rx, 20MB, 5GHz https://github.com/yongsen/SignFi
24 RoGER (Bi et al., 2024) 2 environments, 4 directions, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
25 CSIDA (Zhang et al., 2021c) 2 environments, 5 locations, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/Zhang-xie/WiGr
26 ResMon (Zheng et al., 2023) 2 environments, 3 areas, 6 participants, 4 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
27 Dataset3 (bis, 2024) 2 environments, 6 participants, 4 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
28 FallDeFi (Palipana et al., 2018) 5 environments, 3 participants, fall and others 1Tx, 1Rx, 20MB, 5GHz https://github.com/dmsp123/FallDeFi
29 FallDar (Yang et al., 2022b) 2 environments, 6 locations, 7 participants, fall and normal 1Tx, 1Rx, 20MB, 5GHz http://www.sdp8.net/Dataset?id=dfe4621e-4774-4ef4-b064-e4a082c12335
30 Baha’A et.al. 2020 (Baha’A et al., 2020) 3 environments, 50 participants, 5 actions 1Tx, 1Rx, 20MB, 2.4GHz http://www.sdp8.net/Dataset?id=dfe4621e-4774-4ef4-b064-e4a082c12335
31 SHARP (Meneghello et al., 2022) 3 environments, 4 locations, 3 participants, 4 actions 1Tx, 1Rx, 20MB, 2.4GHz https://github.com/francescamen/SHARP
32 CrossSense (Zhang et al., 2018) 3 environments, 15 locations, 100 participants, 40 actions 1Tx, 1Rx, 20MB, 5GHz Agreement requiredhttps://github.com/nwuzj/CrossSense/
@@
32 GaitID (Zhang et al., 2020c) 2 environment, 7 participants, 16 actions 20MB,5GHz http://www.sdp8.net/Dataset?id=87a65da2-18cb-4b8f-a1ec-c9696890172b
33 NTU-Fi HumanID (Wang et al., 2022e) 1 environment, 20 participants, 6 actions 1Tx, 3Rx,20MB,5GHz https://github.com/xyanchen/WiFi-CSI-Sensing-Benchmark
22 RISE (Zhai et al., 2021) 1 environments, 15 participants, walk 1Tx, 1Rx, 20MB, 5GHz https://github.com/jiaojiao1234/RISE/
30 CrossSense (Zhang et al., 2018) 3 environments, 100 participants, walk 1Tx, 1Rx, 20MB, 5GHz Aggrement required firsthttps://github.com/nwuzj/CrossSense/

### v2

@@ II-D Deployment Stage
II-D1 Transfer Learning with Fine-tuning
@@
The typical steps of transfer learning involve first training a Wi-Fi sensing model on source domain data, followed by updating the model using target domain samples, as shown in Fig. 16(a). This update can be applied to the task head, feature extractor, or the entire model. For example, XRF55 [70], BullyDedect [211], and WiLISensing [153] train models with source domain data and fine-tune the task head using several samples from the target domain, while freezing other parts of the model, enabling few-shot cross-domain activity recognition. DADA-AD [220] fine-tune the deep layers of the model. Besides, Brinke & Meratnia [221] experimentally compare the performance of fine-tuning the task head versus the entire network, finding that fine-tuning the whole network with 20% of target domain data yields the best results. In addition, in some cases, the target domain classification task differs from the source domain, such as in the number of action categories. To address this, WiTransfer [222] replaces the classification head and fine-tunes the new head using target domain data. FewSense [223], on the other hand, discards the classification head and fine-tunes the feature extractor using a cosine similarity loss. This loss pulls features of the same class closer while pushing features of different classes apart. During inference, it employs distance matching for classification, enabling the recognition of novel classes.
@@
Another type of transfer learning consists of two stages: the meta-training stage and the meta-testing stage. In the meta-training stage, a rich labeled dataset from the source domain is used to train an embedding model or feature extractor. In the meta-testing stage, the embedding model is frozen, and a small number of samples from the target domain are input for feature extraction. These features are then used to fine-tune a task head, allowing the model to adapt to the target domain with limited data. Several typical works employ this framework, such as ResMon [224], which uses BayesCNN as the embedding model for domain-adaptive respiration state monitoring. RoGER [225] leverages CNN as the feature extractor to achieve domain-robust gesture recognition. Hou et al. [226] and DASECount [227] both use CNN networks as feature extractors to achieve cross-domain crowd counting.
II-D2 Federated Learning
@@
FDAS [149], WiFederated [228], pFedBKD [229], CARING [230], and Qi et al. [231] utilize federated learning frameworks to achieve efficient model deployment. As shown in Fig. 16(b), the core components include: (1) numerous clients with data collected from diverse environments, (2) a central server that trains a global model with data from diverse environments, which is expected to possess cross-environment generalization capabilities, and (3) the distribution of the global model to client devices, where it is fine-tuned using a small amount of local data. By leveraging these federated learning principles, Wi-Fi sensing models can be effectively deployed in unseen client environments.
@@
Some works adopt frameworks similar to federated learning but use different terminology. For instance, Co-WiSensing [232] proposes a cloud-edge collaborative framework. In this approach, a cloud server trains a cloud Wi-Fi sensing model using data from various edge devices. The higher layers of the trained model are pruned, and the lower layers are distributed to the edge devices. At the edge, the pruned model is extended with a task-specific head, which is fine-tuned using the edge user’s local data to generate a Wi-Fi sensing model tailored for the edge environment. Zhang et al. [119] also employs a central server and adversarial learning to extract shared attributes across different environments. The activity-oriented attributes are then transferred to the local server, enabling cross-environment Wi-Fi human activity recognition. AdaWiFi [233] proposes a collective sensing framework, where a master device collects data using multiple sensors and trains multiple encoders collaboratively, which are then aggregated into a unified model. In the end-user environment, the model is collaboratively tuned with a small amount of labeled data. Zhang et al. [234] propose a local–global modeling method for indoor localization. The approach involves clustering Wi-Fi fingerprints of a large environment into multiple sub-areas, training local localization models for each sub-area, and then aggregating these local models into a global model. The authors demonstrate that this method achieves superior cross-environment generalization.
II-D3 Continual Learning
@@
After model deployment, it is desirable to continuously update the model to adapt to new environments or tasks, as shown in Fig. 16(c). Some works achieve this with online learning [272], a machine learning paradigm where the model learns and updates incrementally as new data becomes available. For instance, FewCS [192] collects a small amount of online data to fine-tune model parameters in new environments, enabling few-shot cross-environment human action recognition. Similarly, WiDrive [235] utilizes online model adaptation and EM algorithm [273] to update the model when in-car driver activities are misclassified, improving recognition for diverse vehicles and drivers. Unlike WiDrive’s automatic error detection, the User-in-the-loop approach allows users to report misclassifications or annotate a small number of samples, which are then used to update the model. M-WiFi [236] specifically asks users to annotate critical time periods to fine-tune the trained model. To minimize the effort required from users, RISE [237] combines probability and statistical assessments with anomaly detection to identify samples that are likely to be misclassified, reducing the need for users to report every error.
@@
Class incremental learning and domain incremental learning are also applied to enable Wi-Fi human sensing models to adapt to new activities or environments. For example, CCS [238] uses a data replay-based incremental learning approach [274], where the model is updated to sense new human activity categories using representative samples from previous stages and samples from the current stage. CAREC [239] is a class-incremental learning framework for Wi-Fi-based indoor action recognition, effectively mitigating catastrophic forgetting [275] through dynamic model expansion and compression, and achieves high accuracy with an 80% reduction in parameters. Similarly, CSI-ARIL [240] treats the environment as a domain, updating the model with representative samples from both old and current environments. CCS, CAREC, and CSI-ARIL ensure that the model does not suffer from catastrophic forgetting when adapting to new classes or domains, preserving its ability to sense old tasks or domains.
III Dataset
Acquiring Wi-Fi sensing data is crucial for effective Wi-Fi sensing applications. Several open-source tools are available for extracting Wi-Fi Channel State Information (CSI) and Beamforming Feedback Information (BFI), as summarized in Table I. Among these, Linux CSI Tool [276] was the earliest open-source CSI extraction tool released in 2011, while PicoScenes currently offers the most advanced CSI extraction capabilities. In addition, due to the limited availability of networks and devices supporting open CSI extraction, BFM-Tool [277] facilitates Beamforming Feedback Information (BFI) extraction, enabling Wi-Fi sensing across a broader range of devices.
TABLE I: Wi-Fi Channel State Information (CSI) and Beamforming Feedback Information (BFI) Extraction Tools
Name
max.
MIMO
802.11
support
max.
# subcarriers
max.
bandwidth
Wi-Fi
Intel CSITool [276] 2$\times$2 n 60 40 MHz CSI
Atheros CSITool [18] - n 114 40 MHz CSI
Nexmon CSITool [278] 4$\times$4 n/ac 256 80 MHz CSI
Wi-ESP [279] 2$\times$2 b/g/n/ac 64 40 MHz CSI
ZTE CSITool [280] 3$\times$2 n/ac/ax 512 160 MHz CSI
PicoScenes [281, 282] 4$\times$4 a/g/n/ac/ax/be 1024 320 MHz CSI
BFM-Tool [277] 4$\times$4 ac/ax 512 160 MHz BFI
Over the years, many researchers have generously released public datasets, significantly improving the efficiency of algorithm evaluation by removing the need to collect data from scratch. Among the various sensing tasks, activity recognition is currently the most well-represented in public datasets, as summarized in Table II. To provide a more intuitive overview of activity recognition datasets, Fig. 17 presents a bubble chart that visualizes the number of activity classes, release date, and dataset volume.
In addition, datasets are also available for tasks such as gait recognition, pose estimation, indoor localization, and temporal action localization. These datasets often include multiple domain attributes, such as environment, participant identity, and action category, making them especially suitable for evaluating generalization and cross-domain performance. In this section, we briefly introduce the domain properties of each dataset to help researchers efficiently identify the most appropriate datasets for evaluating their methods.
[…]
TABLE II: Open-sourced Wi-Fi Human Sensing Datasets
ID Dataset Main Domain Information Device Setting Highlight
@@ III Dataset
1 XRF55 [70] 4 environments, 39 participants, 55 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, 55 actions
2 BullyDetect [211] 8 environments, 20 paired participants, 7 actions 1Tx, 1Rx, 20MHz, 5GHz Bullying actions
3 WiSDA [148] 1 environments, 3 participants, 6 actions 1Tx, 1Rx, 20MHz Widar3.0-like
4 WiNDR [63] 1 environments, 3 participants, 5 actions, 24 orientation 1Tx, 1Rx, 20MHz, 5GHz Full 360° coverage
5 WiMANS [283] 3 environments, 6 participants, 9 actions 1Tx, 1Rx, 20MHz, 2.4/5GHz Multi-person
6 WiGuesture [284] 1 environments, 8 participants, 6 actions 1Tx, 1Rx, 100Hz, 2.4GHz ESP32S3
7 RoGER [225] 2 environments, 4 participants, 6 actions, 4 orientation 1Tx, 2Rx, 20MHz, 2.437GHz AR9580
8 ImgFi [285] 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MHz Converted to images
9 ResMon [224] 2 environments, 3 areas, 6 participants, 4 actions 1Tx, 1Rx, 20/40MHz, 2.4/5GHz Cross-band test
10 Meneghello et al. [286] 6 environments, 4 participants, 7 actions 1Tx, 1Rx, 80MHz, 5.21GHz 80MHz
11 Demrozi et al. [287] 2 environments, 6 participants, 6 actions 2Tx, 2Rx, 2.4GHz Nexmon firmware
12 MM-Fi [72] 4 environments, 40 participants, 27actions 1TX, 1RX, 20MHz, 5GHz Multimodal
13 NTU-Fi HAR [288] 1 environments, 20 participants, 6 actions 1Tx, 1Rx, 40MHz, 5GHz TP-Link N750
14 FallDar [289] 4 environments, 6 locations, 6 participants, fall & normal 1Tx, 1Rx, 1000Hz, 5GHz 6 months
15 ReWis [66] 3 environments, 2 participants, 4 actions 1Tx, 3Rx, 20/80MHz, 5GHz 80MHz
16 OPERAnet [290] 2 environments, 6 participants, 6 actions 1Tx, 2Rx, 5GHz Multimodal
17 SHARP [91] 3 environments, 4 locations, 3 participants, 7 actions 1Tx, 1Rx, 80MB, 5GHz 80MHz
18 CSI-HAR-Dataset [291] 1 environments, 3 participants, 7 actions 1Tx, 1Rx, 20MHz, 5GHz 802.11ac
19 CSIDA [193] 2 environments, 5 locations, 5 participants, 6 actions 1Tx, 1Rx, 40MHz, 5GHz 114 subcarriers
20 RISE [237] 2 environments, 6 participants, 6 actions 1Tx, 1Rx multi-rf devices
21 HTHI [292] 1 environment, 40 paired, 12 actions 2Tx, 3Rx, 20MHz, 2.4GHz Human-human actions
22 DeepSeg [261] 1 environment, 5 participants, 10 actions 1Tx, 3Rx fine-/coarse-grained actions
23 RF-Net [199] 6 environments, 11 participants, 6 actions 2 Tx-Rx pairs, 20MHz 120 environment
24 Baha’A et al. [293] 3 environments, 30 participants, 12 actions 1Tx, 3Rx, 20MHz, 2.4GHz LOS/NLOS
25 Widar 3.0 [67] 3 environments, 5 locations, 5 orientations, 16 participants, 16 actions 1Tx, 6Rx, 20MHz, 5.825GHz Rich domains
26 ARIL [15] 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MHz, 2.4GHz USRP data, clean phase
27 WiAR [294] 3 environments, 10 participants, 16 actions 1Tx, 1Rx, 20MHz, 5GHz RSSI and CSI
28 SAR [295, 221] 1 environment, 9 participants, 6actions 2-3Tx, 3Rx, 20MHz, 2.4GHz 6 days
29 FallDeFi [39] 5 environments, 3 participants, 9 actions 2Tx, 2Rx, 20MHz, 5.2GHz Fall detection
30 SignFi [14] 2 environments, 5 participants, 276 gestures 3Tx, 1Rx, 20MHz, 5GHz 276 ASL Gestures
31 CrossSense [158] 3 environments, 15 locations, 100 participants, 40 actions 1Tx, 1Rx, 20MHz, 5GHz 100 participants
32 Yousef et al. [62] 1 environments, 6 participants, 6 actions 1Tx, 1Rx, 20MHz, 5GHz Fall
@@ 30 SignFi [14] 2 environments, 5 participants, 276 gestures 3Tx, 1Rx, 20MHz, 5GHz 276 ASL Gestures
33 NTU-Fi HumanID [191] 2 environments, 20 participants, walking 1Tx, 1Rx, 40MHz, 5GHz TP-Link N750
20 RISE [237] 3 environments, 15 participants 1Tx, 1Rx multi-rf devices
34 GaitID [30] 2 environments, 11 participants, 8 directions 1Tx, 6Rx 8 directions
31 CrossSense [158] 3 environments, 100 participants 1Tx, 1Rx, 20MHz, 5GHz 100 participants